The recovery rate from a flu is 0.9. If 4 people have this flu, what is the probability that:
a) all will recover?
b) at least 2 will recover?
c) exactly 2 will recover?
Assume independence.
Thanks so much for your help!

Answer by stanbon(75887)

The recovery rate from a flu is 0.9. If 4 people have this flu, what is the probability that:
a) all will recover?
Ans: (0.9)^4 =
--------------------------
b) at least 2 will recover?
Ans: P(2<= x <=4)
= 1 - P(0<= x <=2)
= 1-binomcdf(4,0.9,2)
= 1-0.0523
= 0.9477
====================
c) exactly 2 will recover?
Ans: P(x = 2) = 4C2(0.9)^2(0.1)^2 = binompdf(4,0.9,2) = 0.048600
====================================================================
Cheers,
Stan H.
================